# Perturbative corrections to soft photon theorems for massless scalar QED in de Sitter spacetime

### v1

@@ 2.1 De Sitter spacetime
We study the scattering of massless scalars followed by an emission of a graviton in the static patch of the de Sitter space. We confine this scattering process to the small compact region $R$ inside the static patch as shown in Fig. (1). The de Sitter metric can be put in the conformally flat form in the stereographic coordinates $x^{\mu}$ as is shown in [26, 25], i.e,
@@ 2.2 Key assumptions in the setup
    where $\Gamma_{n}$ is the $\mathcal{S}$-matrix for the scattering of massless $n$-scalars without an emission of the photon, $A^{(0)}$ is the Weinberg soft factor, and $A^{(1)}$ and $A^{(2)}$ are the de Sitter corrections. These perturbative corrections are only valid for $\omega l>>1$. Also, the soft limit in the flat space requires $\omega<<E$, where $E$ is the energy of the hard particles, i.e., scalars, that participate in the scattering process. Thus, in the present scenario, the soft limit is valid in the regime of $\frac{1}{l}<<\omega<<E$. Since $l$ is very large and $\omega\to 0$, it is convenient to introduce a constant parameter $\delta=\omega l$. In this new parameter, the total soft factor can be expanded as follows:
@@ 4.1 Scalar modes
We can express the creation and annihilation operators appearing above in terms of field operators by using the inner product defined in (22). This deduction has been done in [25] and we quote the result here. For the annihilation operator corresponding to the scalar field, we have
@@ 5 Perturbative $\mathcal{S}$-matrix and soft factor corrections
In this section, we define a perturbative $\mathcal{S}$-matrix in the de Sitter space and deduce curved spacetime generalization of the LSZ formula. The LSZ formula in [26] was obtained for massive scalars. Here, we focus on massless scalars. Thus, in a sense, the perturbative $\mathcal{S}$-matrix for the present case should be obtained in a suitable limit of the scalar mass, $m\rightarrow 0$. Another point to mention is that to formulate a convenient $\mathcal{S}$-matrix, we need to prescibe asymptotic states at the boundaries of spacetime. However, in de Sitter space, there is no notion of a global observable boundary and the only observable region is the static patch. This patch is bounded by the cosmological horizons. In this situation, we defer from defining proper asymptotic states at null infinity but instead treat the cosmological horizons as physical boundaries of this spacetime  [^1] In [31, 32], the cosmological horizons are considered as physical boundaries to study the thermodynamics of a rotating black hole. to define the so called asymptotic states. In our analysis, since the scattering processes are confined to the compact region inside the static patch, the asymptotic states are defined at the early- and late-time Cauchy surfaces which act as Hilbert spaces of the incoming and outgoing particles. The particles can reach the Cauchy surface in a finite time $T$, which we assume to be much larger than the interaction time scale. All lengths in our calculations are very small compared to the de Sitter radius $l$, i.e., $|x|<<l$.
The $\mathcal{S}$-matrix describing the scattering of incoming particles with momenta $p_{i}$ to the outgoing particles with momenta $p_{j}$ followed by the emission of gravitons with momenta $p_{k}$ is defined as
@@ 5.1 Soft photon scattering
$\Gamma_{n+1}(\{p_{i}\},k)=S(p_{i},k)\Gamma_{n}(\{p_{i}\}),$ (33)
where $p_{i}$ are the momenta of hard particles, $k$ is the soft momentum and $S(p_{i},k)$ is the soft factor.
Due to the presence of de Sitter background, the flat space soft factor as well as the amplitude $\Gamma_{n}(\{p_{i}\})$ receives $O(\frac{1}{\ell})$ corrections.
@@
where $G[\phi(z)...\phi(z_{j})]$ is the $n$-point correlation function which receives contributions from any arbitrary interaction vertices, $M_{n+1}$ is the $\mathcal{S}$-matrix for the second diagram in Fig. (2) where the graviton is not attached to an external line but to an internal line. The summation has been taken over all the scalars since any of them can emit the photon. This is the main result from which we compute the leading, subleading, and sub-subleading corrections to the flat space soft factor in the following sections. Substituting all the essential elements, i.e., scalar field modes, graviton modes, and the scalar field propagator, in the above expression, keeping terms up to $\mathcal{O}(l^{-2})$, and taking the soft limit ($k\to 0$ or $\omega\to 0$) produces the final expression that can be put in the following form:
$\Gamma_{n+1}(\{p_{1},...,p_{n}\},\omega\hat{k})=\bigg(A^{(0)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l}A^{(1)}(\{p_{i}\},\omega\hat{k})+\dfrac{1}{l^{2}}A^{(2)}(\{p_{i}\},\omega\hat{k})\bigg)\\ \cdot\bigg(\Gamma_{n}^{(0)}(\{p_{1},...,p_{n}\})+\dfrac{1}{l}\Gamma_{n}^{(1)}(\{p_{1},...,p_{n}\})+\dfrac{1}{l^{2}}\Gamma_{n}^{(2)}(\{p_{1},...,p_{n}\})\bigg),$ (35)
where $A^{(1)}$ and $A^{(2)}$ are the perturbative corrections to the flat space soft factor $A^{(0)}$. Similarly, $\Gamma_{n}^{(1)}$ and $\Gamma_{n}^{(2)}$ are the perturbative corrections to the flat space $\mathcal{S}$-matrix, $\Gamma_{n}^{(0)}$, for the scattering of $n$-scalars.
We must note that the perturbative corrections to the flat space soft factor are only valid for $\omega l>>1$. And the soft limit of scattering amplitudes in flat space requires $\omega<<p_{i}^{\mu}$, where $p_{i}^{\mu}$ is the momenta of the hard particles. Thus, our analysis, and hence the results are valid in the regime $\frac{1}{l}<<\omega<<p_{i}^{\mu}$. For the soft graviton scattering in the de Sitter background, we take the double scaling limit [23] in which $\omega\to 0$ and $l\to\infty$ simultaneously, keeping their product $\omega l$ finite. In these limits, it is convenient to introduce a new parameter $\delta=1/\omega l$. To put the perturbative corrections in a proper hierarchy of leading, subleading, and sub-subleading corrections, the full soft factor (i.e. flat space + perturbative corrections) can be expanded in $\delta$ as follows:
[…]
where ‘$s$’ represents the power of $\omega$, i.e., $S^{(r,s)}\sim\mathcal{O}(\omega^{s})$. In this soft expansion, $S^{(0,-1)}$ is the flat space leading soft factor, and the higher order terms are leading, subleading, sub-subleading corrections, and so on.
@@ Acknowledgements
PC thanks Divyesh N Solanki for various conversations about this work. PC is supported in part by the National Natural Science Foundation of China (Grant No. 12475105).

### v2

@@ 2.1 De Sitter spacetime
We study the scattering of massless scalars followed by an emission of a photon in the static patch of the de Sitter space. We confine this scattering process to the small compact region $R$ inside the static patch as shown in Fig. (1). The de Sitter metric can be put in the conformally flat form in the stereographic coordinates $x^{\mu}$ as is shown in [26, 25], i.e,
@@ 2.2 Key assumptions in the setup
    where $\Gamma_{n}$ is the $\mathcal{S}$-matrix for the scattering of massless $n$-scalars without an emission of the photon, $A^{(\text{flat})}$ is the Weinberg soft factor, and $A^{(\text{dS})}$ is the de Sitter correction. These perturbative corrections are only valid for $\omega l>>1$. Also, the soft limit in the flat space requires $\omega<<E$, where $E$ is the energy of the hard particles, i.e., scalars, that participate in the scattering process. Thus, in the present scenario, the soft limit is valid in the regime of $\frac{1}{l}<<\omega<<E$. Since $l$ is very large and $\omega\to 0$, it is convenient to introduce a constant parameter $\delta=\omega l$. In this new parameter, the total soft factor can be expanded as follows:
@@ 4.1 Scalar modes
We can express the creation and annihilation operators appearing above in terms of field operators by using the inner product defined in (22). This deduction has been done in [25] but in their case, it seems they have considered a neutral scalar field as their mode solution is real. This is not legitimate and hence we take a complex mode solution. Now the important point we is that the LSZ and the S matrix does not change if some of the scalars are anti particles instead of particles. This is due to crossing symmetry. Therefore, where confusions do not arise, we consider all scalars to be particles (and not antiparticles or a mixture). We then return to their result and quote it here. For the annihilation operator corresponding to the scalar field, we have
@@ 5 Perturbative $\mathcal{S}$-matrix and soft factor corrections
In this section, we define a perturbative $\mathcal{S}$-matrix in the de Sitter space and deduce curved spacetime generalization of the LSZ formula. The LSZ formula in [26] was obtained for massive scalars. Here, we focus on massless scalars. Thus, in a sense, the perturbative $\mathcal{S}$-matrix for the present case should be obtained in a suitable limit of the scalar mass, $m\rightarrow 0$. Another point to mention is that to formulate a convenient $\mathcal{S}$-matrix, we need to prescribe asymptotic states at the boundaries of spacetime. However, in de Sitter space, there is no notion of a global observable boundary and the only observable region is the static patch. This patch is bounded by the cosmological horizons. In this situation, we defer from defining proper asymptotic states at null infinity but instead treat the cosmological horizons as physical boundaries of this spacetime  [^1] In [31, 32], the cosmological horizons are considered as physical boundaries to study the thermodynamics of a rotating black hole. to define the so called asymptotic states. In our analysis, since the scattering processes are confined to the compact region inside the static patch, the asymptotic states are defined at the early- and late-time Cauchy surfaces which act as Hilbert spaces of the incoming and outgoing particles. The particles can reach the Cauchy surface in a finite time $T$, which we assume to be much larger than the interaction time scale. All lengths in our calculations are very small compared to the de Sitter radius $l$, i.e., $|x|<<l$.
The $\mathcal{S}$-matrix describing the scattering of incoming particles with momenta $p_{i}$ to the outgoing particles with momenta $p_{j}$ followed by the emission of photons with momenta $p_{k}$ is defined as
@@ 5.1 Soft photon scattering
$\Gamma_{n+1}(\{p_{i}\},k)=A(\{p_{i}\},k)\Gamma_{n}(\{p_{i}\}),$ (33)
where $p_{i}$ are the momenta of hard particles, $k$ is the soft momentum and $A(\{p_{i}\},k)$ is the soft factor.
Due to the presence of de Sitter background, the flat space soft factor as well as the amplitude $\Gamma_{n}(\{p_{i}\})$ receives $O(\frac{1}{\ell^{2}})$ corrections.
@@
where $G[\phi(z)...\phi(z_{j})]$ is the $n$-point correlation function which receives contributions from any arbitrary interaction vertices, $M_{n+1}$ is the $\mathcal{S}$-matrix for the second diagram in Fig. (2) where the photon is not attached to an external line but to an internal line. The summation has been taken over all the scalars since any of them can emit the photon. This is the main result from which we compute the leading, subleading, and sub-subleading corrections to the flat space soft factor in the following sections.
@@ Acknowledgements
PC thanks Divyesh N Solanki for suggesting corrections on a version of this draft and for various conversations about this work. PC is supported in part by the National Natural Science Foundation of China (Grant No. 12475105).